OpenStudy (anonymous):

What is the slope of the line represented? Y-3=-2(x+1)

OpenStudy (tyteen4a03):

Expand -2(x+1), then solve the equation until you get the slope-intercept formula of a line: y = mx + c, where m is the slope.

OpenStudy (anonymous):

Great! Thanks!

OpenStudy (anonymous):

Easier way. This equation is in point-slope form. (y-y1) = m (x-x1). The slope is the m.